# Multi-messenger lensing time delay as a probe of the graviton mass

### v1

@@ Abstract
Gravitational lensing is a powerful probe of cosmology and astrophysics. With the prospect of the first strongly lensed gravitational waves on the horizon, we highlight an opportunity to test fundamental physics. In this work, we assume a nonzero mass for the graviton, which leads to gravitational waves following timelike geodesics instead of null geodesics. We derive standard gravitational lensing equations, such as the scattering angle, the time-delay between different images and the magnification, which normally rely on the assumption of null geodesics. We show that a single strongly lensed multi-messenger event is enough to constrain the graviton mass to $m<3\cdot 10^{-23}$ eV/c2. Notably this constraint is independent of the lens model, the waveform model, and of cosmology. Additionally, we explore magnification of images and find that they offer at least three orders of magnitude weaker bounds than the time delay, and have a dependence on the correct modeling of the lens and cosmology.
@@ I Introduction
Strong gravitational lensing of electromagnetic waves is a well-established probe of cosmology [1]. In recent years gravitational wave (GW) strong lensing has been actively developing both theoretically and observationally [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] as a complimentary method to EM lensing – though searches have yet to find conclusive lensed signals. This is due to the rarity of these events combined with current detector sensitivity limitations [13]; we expect to observe one lensed event for every $\sim 1500$ unlensed detections [11]. [^1] To contextualise this, approximately 200 unlensed events have been reported so far. The combination of EM and GW gravitational lensing naturally leads to the possibility of multi-messenger lensing [14]: a multi-messenger strongly lensed event, which is often called a ‘golden’ scenario. Despite the predicted scarcity of strongly lensed GW signals, if one is observed it will be straightforward to find the lensed host galaxy, thanks to new wide-field observatories such as Euclid [15] and LSST [16]. Prospects are quite optimistic due to the sheer number of strongly lensed galaxies: the recent first data release from the Euclid mission [17] reported 497 new strong lenses, and forecasts predict a total of $\sim 75,000$ lenses by the end of the survey [18]. Additionally, LSST forecasts $\sim 70,000$ strong lenses (for a conservative estimate) in 10 years [19], with more optimistic predictions up to order $\sim 400,000$ strong lenses from Euclid, LSST, and DES combined [20]. Detection of a golden event would offer the opportunity to perform tests of general relativity in addition to cosmological constraints.
@@
Modified gravity models have been explored in the context of unlensed multi-messenger signals [21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. In the context of gravitational wave lensing, a few studies of modified gravity have been performed. For example, different authors investigated tests of local Lorentz violations [31], anomalous GW damping [32, 33] anomalous speed of GWs [34, 35] or both [36]. The effect of Horndeski gravity on a lens was studied in [37] and gravitational lensing of electromagnetic signals was investigated for a subset of Horndeski theories in which GWs travel at light speed in [38]. The frequency-dependent amplification factor of lensed gravitational waves was proposed as a novel test of the graviton mass in [39].
@@
In this work, we bridge the gap between lensing in massive gravity and multi-messenger gravitational wave events. We perform a comprehensive theoretical calculation for a golden event in the case of a massive graviton. We investigate how the presence of a mass in the dispersion relation of gravitational waves affects geodesics, time delays, and magnification in a gravitational lensing context. We outline observational prospects of a golden event in this scenario. We show how comparing EM and GW signals allows to impose model-independent bounds on the graviton mass, adding to the existing dispersion relation tests of gravity which find $m\leq 1.27\cdot 10^{-23}$ eV/c2 (90% confidence level) [40].
@@
We show that a fully model-independent bound on the graviton mass may be imposed from the time delay of a strongly lensed multi-messenger event, which will provide an alternative constraint, complementary to both GW dispersion constraints mentioned above and non-GW bounds from the Solar System, clusters, and weak lensing (see [41] for a comprehensive review).
@@
This article is structured as follows. In Sec. II, we show how a mass may affect the dispersion relation of gravitational waves and derive the geodesic equation. In Sec. III, we show how the scattering angle of a lensed GW is affected by the mass of the graviton. In Sec. IV, we compute the time delay between different images for a strongly lensed massive gravitational wave and explore multi-messenger constraints. In Sec. V, we derive the magnification of massive GWs from scratch and show that the constraints that can be obtained on the graviton mass from the comparison of EM versus GW magnification is weaker than from the time delay. Finally, we conclude in Sec. VI. We chose the (,+,+,+)(-,+,+,+) metric signature and units are such that $c=1=\hbar$.
@@ II Massive geodesics
In this section, we make a pedagogical introduction on how a mass term introduced in the field equations may alter the geodesics of gravitational waves. We study this phenomenologically and defer more formal aspects of massive gravity to the literature [42] and references therein. In this context, a modification of the dispersion relation can arise if the linearized vacuum Einstein field equations in the Hilbert gauge ($\nabla^{\mu}h_{\mu\nu}=0$) admit an extra mass term
$\Box h_{\mu\nu}-2h^{\alpha\beta}R_{\alpha\mu\nu\beta}-m^{2}h_{\mu\nu}=0\,$, (1)
where we split the metric $\bar{g}_{\mu\nu}=g_{\mu\nu}+h_{\mu\nu}$ into a background spacetime described by $g_{\mu\nu}$ and gravitational waves, described by $h_{\mu\nu}$. In Eq. (1), the Riemann tensor $R_{\alpha\mu\nu\beta}$, which arises in a generically curved background, and $\Box\equiv g^{\mu\nu}\nabla_{\mu}\nabla_{\nu}$ operator are defined with respect to the background metric $g_{\mu\nu}$. We make the following wave Ansatz
@@
where $H_{\mu\nu}$ describes the real amplitude and polarization of the wave, while $\varphi$ describes the phase. The wavevector is defined as $k_{\mu}\equiv\nabla_{\mu}\varphi$. Plugging Eq. (2) into Eq. (1), we get
$-(k^{\alpha}k_{\alpha}+m^{2})H_{\mu\nu}$ $+\left(\Box H_{\mu\nu}-2H^{\alpha\beta}R_{\alpha\mu\nu\beta}\right)$ (3)
@@
which describes the dispersion relation of gravitational waves. While we have added the mass term by hand in the linearized field equations for pedagogical purposes, we consider this dispersion relation Eq. (5) to be the starting assumption of this work. The imaginary part (Eq. 4) can be integrated to solve for the amplitude of the gravitational wave and to show that the polarization is parallel transported along the geodesic to leading order in geometric optics (see for example [45]).
@@
$0$ $=k^{\mu}(\nabla_{\nu}k_{\mu})=k^{\mu}(\partial_{\nu}\partial_{\mu}\varphi-\Gamma^{\lambda}_{\nu\mu}k_{\lambda})$ (6)
$=k^{\mu}(\partial_{\mu}\partial_{\nu}\varphi-\Gamma^{\lambda}_{\mu\nu}k_{\lambda})$ (7)
$=k^{\mu}\nabla_{\mu}k_{\nu}$ (8)
@@
where subscripts $o$ and $s$ refer to the position of the observer and the source, respectively. From this expression, it seems that given the same initial direction of propagation, a massive and massless graviton will generally travel along different curves since the background Christoffel symbols contracts with all components of $k_{\mu}$, which satisfy different constraints: the massive case follows Eq. (5), while in the case of a massless graviton, we would have $k^{\mu}k_{\mu}=0$. In the next section, we illustrate the different scattering angles for massive and massless gravitons in a simple setting of a point-like lens.
@@ III Scattering angle
One can easily check that $k^{\mu}k_{\mu}=-m^{2}$ with $\omega^{2}=k^{2}+m^{2}$. The meaning of the affine parameter can be understood by taking the projection $d_{\mu}$ on a infinitesimal coordinate displacement $\mathrm{d}x^{\mu}$. In the frame of the observer, the distance traveled by a photon corresponds to
@@
Choosing an orientation, this wavevector can be written as $\bar{k}^{\mu}(\lambda)=(\omega,0,0,k)$, which describes the geodesic of a massive graviton traveling in Minkowski space towards the $\boldsymbol{\hat{e}_{z}}$ direction. If the wave is sent from $\boldsymbol{x}=(x=b,y=0,z\to-\infty)$, where $b$ is the impact parameter, we expect the wave to be deflected towards the $\boldsymbol{\hat{e}_{x}}$ direction.
For a wave propagating towards the $\boldsymbol{\hat{e}_{z}}$ direction, this implies that $\mathrm{d}z=k\mathrm{d}\lambda$, as per Eq. (15). We now compute the integrals which describe $\delta k^{i}$. To first order in $U\ll 1$, Eq. (9) leads to
$\delta k^{i}(\lambda_{o})=-\int_{\mathbb{R}}\frac{\mathrm{d}z}{k}\left(\Gamma^{i}_{00}(\bar{k}^{0})^{2}+\Gamma^{i}_{jk}\bar{k}^{j}\bar{k}^{k}\right)$ (16)
@@
Therefore, the normalised deflected wavevector $\boldsymbol{\hat{k}}$ at the observer, which we denote with a hat, reads
@@
where the $m^{2}/\omega^{2}$ holds the leading order difference between the deflection angle of a massive and massless geodesic. Massive gravitons are more deflected than massless particles. Their images form at
@@ IV Time-delay constraint
In this section, we start from the massive geodesic equation and derive the time delay between two images strongly lensed by a foreground lens in the context of a FLRW background spacetime. We first compute the geometrical time delay in Sec. IV.1 then focus on the Shapiro time delay in Sec. IV.2. The total time delay is the sum of these two contributions, as detailed in Sec. IV.3.
@@ IV.1 Geometric time delay
We consider a flat FLRW metric in conformal time, which is described by the following line element,
@@
The first integral represents the comoving distance between $\lambda_{1}$ and $\lambda_{2}$, while the second integral is a mass correction term. The geometrical time delay between two signals emitted simultaneously and traveling the deflected versus the undeflected path read
@@
Next, we can use the relation between comoving distance and angular diameter distance $D_{s}=d_{s}/(1+z_{s})$, $D_{l}=d_{l}/(1+z_{l})$, $D_{ls}=d_{ls}/(1+z_{s})$ and $D_{\ell s}\boldsymbol{\hat{\alpha}}=D_{s}(\boldsymbol{\theta}-\boldsymbol{\beta})$, which implies that we can rewrite the geometric time delay as
@@
where $z_{l}$ is the redshift of the lens, $D_{s}$, $D_{l}$ and $D_{ls}$ are the angular diameter distances to the source, to the lens and between the lens and the source, while $\boldsymbol{\theta}$ and $\boldsymbol{\beta}$ are the angular positions of the image and source respectively. The formula also holds for a non-flat FLRW geometry [47].
@@ IV.2 Shapiro time delay
In this section, we compute the second contribution of the time delay, the Shapiro time delay. For this purpose, we assume that the metric is well described by the line element in Eq. (10), with the static weak field gravitational potential $U(\boldsymbol{x})\ll 1$. Note that we use a different line element for this contribution, as we take into account the effect of the gravitational potential of the lens, for which cosmology is irrelevant. We rearrange Eq. (5) into
$k^{0}\equiv\frac{\mathrm{d}t}{\mathrm{d}\lambda}$ $=\pm\sqrt{-\frac{g_{ij}}{g_{00}}\frac{\mathrm{d}x^{i}}{\mathrm{d}\lambda}\frac{\mathrm{d}x^{j}}{\mathrm{d}\lambda}-\frac{m^{2}}{g_{00}}}$ (30)
@@
where the first term in the second bracket is the path length along the affine parameter, the contribution to the time delay of which we have calculated in detail in the previous section (see Sec. IV.1, Eq. 29). The second term in the second bracket is the Shapiro time delay, which arises because of the presence of the gravitational field [47]. This contribution depends on the path $\gamma(\lambda)$. The graviton mass acts as a correction to the Shapiro time delay.

### v2

# Multimessenger lensing time delay as a probe of the graviton mass

@@ Abstract
Gravitational lensing is a powerful probe of cosmology and astrophysics. With the prospect of the first strongly lensed gravitational waves on the horizon, we highlight an opportunity to test fundamental physics. In this work, we assume a nonzero mass for the graviton, which leads to gravitational waves following timelike geodesics instead of null geodesics. We derive standard gravitational lensing equations, such as the scattering angle, the time delay between different images and the magnification, which normally rely on the assumption of null geodesics. We show that a single strongly lensed multimessenger event is enough to constrain the graviton mass to $m<3\cdot 10^{-23}$eV/c2. Notably this constraint is independent of the lens model, of the waveform model, and of cosmology. Additionally, we explore magnification of images and find that they offer at least 3 orders of magnitude weaker bounds than the time delay, and have a dependence on the correct modeling of the lens and cosmology.
@@ I Introduction
Strong gravitational lensing of electromagnetic (EM) waves is a well-established probe of cosmology [16]. In recent years gravitational wave (GW) strong lensing has been actively developing both theoretically and observationally [68, 65, 63, 25, 49, 41, 34, 35, 20, 69, 67, 3] as a complementary method to EM lensing – though searches have yet to find conclusive lensed signals. This is due to the rarity of these events combined with current detector sensitivity limitations [3]; we expect to observe one lensed event for every $\sim 1500$ unlensed detections [69]. [^1] To contextualize this, approximately 200 unlensed events have been reported so far. The combination of EM and GW gravitational lensing naturally leads to the possibility of multimessenger lensing [64]: a multimessenger strongly lensed event, which is often called a “golden” scenario. Despite the predicted scarcity of strongly lensed GW signals, if one is observed it will be straightforward to find the lensed host galaxy, thanks to new wide-field observatories such as Euclid [54] and LSST [44]. Prospects are quite optimistic due to the sheer number of strongly lensed galaxies: the recent first data release from the Euclid mission [5] reported 497 new strong lenses, and forecasts predict a total of $\sim 75,000$ lenses by the end of the survey [50]. Additionally, LSST forecasts $\sim 70,000$ strong lenses (for a conservative estimate) in 10 years [61], with more optimistic predictions up to order $\sim 400,000$ strong lenses from Euclid, LSST, and DES combined [24]. Detection of a golden event would offer the opportunity to perform tests of general relativity in addition to cosmological constraints.
@@
Modified gravity models have been explored in the context of unlensed multimessenger signals [4, 9, 8, 47, 6, 28, 29, 7, 22, 36]. In the context of gravitational wave lensing, a few studies of modified gravity have been performed. For example, different authors investigated tests of local Lorentz violations [14], anomalous GW damping [39, 57] anomalous speed of GWs [23, 38] or both [37]. The effect of Horndeski gravity on a lens was studied in [46] and gravitational lensing of electromagnetic signals was investigated for a subset of Horndeski theories in which GWs travel at light speed in [13]. The frequency-dependent amplification factor of lensed gravitational waves was proposed as a novel test of the graviton mass in [21].
@@
In this work, we bridge the gap between lensing in massive gravity and multimessenger gravitational wave events. We perform a comprehensive theoretical calculation for a golden event in the case of a massive graviton. We investigate how the presence of a mass in the dispersion relation of gravitational waves affects geodesics, time delays, and magnification in a gravitational lensing context. We outline observational prospects of a golden event in this scenario. We show how comparing EM and GW signals allows one to impose model-independent bounds on the graviton mass, adding to the existing dispersion relation tests of gravity, which find $m\leq 1.27\cdot 10^{-23}$eV/c2 (90% confidence level) [2].
@@
We show that a fully model-independent bound on the graviton mass may be imposed from the time delay of a strongly lensed multimessenger event, which will provide an alternative constraint, complementary to both GW dispersion constraints mentioned above and non-GW bounds from the Solar System, clusters, binary pulsars [58] and weak lensing (see [30] for a comprehensive review).
@@
This article is structured as follows. In Sec. II, we show how a mass may affect the dispersion relation of gravitational waves and derive the geodesic equation. In Sec. III, we show how the scattering angle of a lensed GW is affected by the mass of the graviton. In Sec. IV, we compute the time delay between different images for a strongly lensed massive gravitational wave and explore multimessenger constraints. In Sec. V, we derive the magnification of massive GWs from scratch and show that the constraints that can be obtained on the graviton mass from the comparison of EM versus GW magnification is weaker than from the time delay. Finally, we conclude in Sec. VI. We chose the (,+,+,+)(-,+,+,+) metric signature and units are such that $c=1=\hbar$.
@@ II Massive geodesics
In this section, we make a pedagogical introduction on how a mass term may alter the geodesics of gravitational waves. We study this phenomenologically and defer more formal aspects of massive gravity to the literature [32] and references therein. Note that a UV completion of Lorentz violating massive gravity was proposed in [17]. Suppose that the equation of motion for the metric perturbation reads
$(\Box-m^{2})h_{\mu\nu}+2R_{\mu\rho\nu\sigma}h^{\rho\sigma}\simeq 0\,$, (1)
where we split the metric $\bar{g}_{\mu\nu}=g_{\mu\nu}+h_{\mu\nu}$ into a background spacetime described by $g_{\mu\nu}$ and gravitational waves, described by $h_{\mu\nu}$. In Eq. (1), the box operator and the Riemann tensor $R_{\mu\rho\nu\sigma}$ are defined with respect to the background metric $g_{\mu\nu}$, i.e. $\Box\equiv g^{\mu\nu}\nabla_{\mu}\nabla_{\nu}$ and $m$ denotes the graviton mass. Such an equation of motion may arise from the variation of the following action for a massive graviton (see for example [10])
$S[h_{\mu\nu}]$ $=-M_{\rm Pl}^{2}\int\mathrm{d}^{4}x\sqrt{-g}\Big[h_{\mu\nu}\mathcal{D}^{\mu\nu\rho\sigma}h_{\rho\sigma}$ (2)
$\quad-\Lambda\left(h_{\mu\nu}h^{\mu\nu}-\frac{1}{2}h^{2}\right)+\frac{m^{2}}{2}\left(h_{\mu\nu}h^{\mu\nu}-\frac{1}{2}h^{2}\right)\Big]$
where $M_{\rm Pl}$ denotes the Planck mass, $\Lambda$ denotes the cosmological constant, $h=g^{\mu\nu}h_{\mu\nu}$ and the operator
$\mathcal{D}^{\mu\nu\rho\sigma}h_{\rho\sigma}$ $\equiv-\frac{1}{2}\Big[g^{\mu\rho}g^{\nu\sigma}\Box+g^{\rho\sigma}\nabla^{\mu}\nabla^{\nu}-g^{\mu\rho}\nabla^{\sigma}\nabla^{\nu}$ (3)
$\quad-g^{\nu\rho}\nabla^{\sigma}\nabla^{\mu}-g^{\mu\nu}g^{\rho\sigma}\Box+g^{\mu\nu}\nabla^{\rho}\nabla^{\sigma}\Big]h_{\rho\sigma}\,.$
The equation of motion associated with the action in (2), Eq. (1) follows with the constraints $\nabla^{\mu}h_{\mu\nu}\simeq 0$, $h\simeq 0$, if the background spacetime satisfies $R_{\mu\nu}=\Lambda g_{\mu\nu}$. While there exists this formal possibility, the relevant aspect for our work is the fact that the metric perturbation factors out of the $(\Box-m^{2})$ term in Eq. (1). Note that a more complicated structure arises for de Rham-Gabadadze-Tolley (dRGT) [31] ghost-free massive gravities [12, 11].
Adopting Eq. (1) as the equation of motion, we make the following wave Ansatz
@@
where $H_{\mu\nu}$ describes the real amplitude and polarization of the wave, while $\varphi$ describes the phase. In principle, $H_{\mu\nu}$ can describe up to five polarization modes, which follows from the 10 degrees of freedom of a symmetric rank 2 tensor deducted from the 5 constraints $\nabla^{\mu}h_{\mu\nu}\simeq 0\simeq h$. The wave vector is defined as $k_{\mu}\equiv\nabla_{\mu}\varphi$. Plugging Eq. (4) into Eq. (1), we get
$-(k^{\alpha}k_{\alpha}+m^{2})H_{\mu\nu}$ $+\left(\Box H_{\mu\nu}-2H^{\alpha\beta}R_{\mu\alpha\nu\beta}\right)$
@@
which describes the dispersion relation of gravitational waves. While we have presented an example action that leads to Eq. (6) for pedagogical purposes, we consider this dispersion relation Eq. (6) to be the starting assumption of this work. Since we have chosen the phase $\varphi$ to be polarization independent, Eq. (6) applies to any polarization. The imaginary part [Eq. (5)] can be integrated to solve for the amplitude and polarization of the gravitational wave and to show that the polarization is parallel transported along the geodesic to leading order in geometric optics (see for example [27]).
@@
$0$ $=k^{\mu}(\nabla_{\nu}k_{\mu})=k^{\mu}(\partial_{\nu}\partial_{\mu}\varphi-\Gamma^{\lambda}_{\nu\mu}k_{\lambda})$
$=k^{\mu}(\partial_{\mu}\partial_{\nu}\varphi-\Gamma^{\lambda}_{\mu\nu}k_{\lambda})$
$=k^{\mu}\nabla_{\mu}k_{\nu}\,$, (7)
@@
where subscripts $o$ and $s$ refer to the position of the observer and the source, respectively. From this expression, it seems that given the same initial direction of propagation, a massive and massless graviton will generally travel along different curves since the background Christoffel symbols contracts with all components of $k_{\mu}$, which satisfy different constraints: the massive case follows Eq. (6), while in the case of a massless graviton, we would have $k^{\mu}k_{\mu}=0$. In the next section, we illustrate the different scattering angles for massive and massless gravitons in a simple setting of a pointlike lens.
@@ III Scattering angle
One can easily check that $k^{\mu}k_{\mu}=-m^{2}$ with $\omega^{2}=k^{2}+m^{2}$. The meaning of the affine parameter can be understood by taking the projection $d_{\mu}$ on an infinitesimal coordinate displacement $\mathrm{d}x^{\mu}$. In the frame of the observer, the distance traveled by a photon corresponds to
@@
Choosing an orientation, this wave vector can be written as $\bar{k}^{\mu}(\lambda)=(\omega,0,0,k)$, which describes the geodesic of a massive graviton traveling in Minkowski space toward the $\boldsymbol{\hat{e}_{z}}$ direction. If the wave is sent from $\boldsymbol{x}=(x=b,y=0,z\to-\infty)$, where $b$ is the impact parameter, we expect the wave to be deflected toward the $\boldsymbol{\hat{e}_{x}}$ direction.
For a wave propagating toward the $\boldsymbol{\hat{e}_{z}}$ direction, this implies that $\mathrm{d}z=k\mathrm{d}\lambda$, as per Eq. (14). We now compute the integrals which describe $\delta k^{i}$. To first order in $U\ll 1$, Eq. (8) leads to
$\delta k^{i}(\lambda_{o})=-\int_{\mathbb{R}}\frac{\mathrm{d}z}{k}\left(\Gamma^{i}_{00}(\bar{k}^{0})^{2}+\Gamma^{i}_{jk}\bar{k}^{j}\bar{k}^{k}\right)\,.$ (15)
@@
Therefore, the normalized deflected wave vector $\boldsymbol{\hat{k}}$ at the observer, which we denote with a hat, reads
@@
where the $m^{2}/\omega^{2}$ holds the leading order difference between the deflection angle of a massive and a massless geodesic. Massive gravitons are more deflected than massless particles. Their images form at
@@ IV Time-delay constraint
In this section, we start from the massive geodesic equation and derive the time delay between two images strongly lensed by a foreground lens in the context of a FLRW (Friedmann–Lemaître–Robertson–Walker) background spacetime. We first compute the geometrical time delay in Sec. IV.1 and then focus on the Shapiro time delay in Sec. IV.2. The total time delay is the sum of these two contributions, as detailed in Sec. IV.3.
@@ IV.1 Geometric time delay
We consider a flat FLRW metric in conformal time, which is described by the following line element:
@@
The first integral represents the comoving distance between $\lambda_{1}$ and $\lambda_{2}$, while the second integral is a mass correction term. The geometrical time delay between two signals emitted simultaneously and traveling the deflected versus the undeflected path reads
@@
Next, we can use the relation between the comoving distance and the angular diameter distance $D_{s}=d_{s}/(1+z_{s})$, $D_{l}=d_{l}/(1+z_{l})$, $D_{ls}=d_{ls}/(1+z_{s})$ and $D_{\ell s}\boldsymbol{\hat{\alpha}}=D_{s}(\boldsymbol{\theta}-\boldsymbol{\beta})$, which implies that we can rewrite the geometric time delay as
@@
where $z_{l}$ is the redshift of the lens, $D_{s}$, $D_{l}$ and $D_{ls}$ are the angular diameter distances to the source, to the lens and between the lens and the source, while $\boldsymbol{\theta}$ and $\boldsymbol{\beta}$ are the angular positions of the image and source respectively. The formula also holds for a nonflat FLRW geometry [59].
@@ IV.2 Shapiro time delay
In this section, we compute the second contribution of the time delay, the Shapiro time delay. For this purpose, we assume that the metric is well described by the line element in Eq. (9), with the static weak-field gravitational potential $U(\boldsymbol{x})\ll 1$. Note that we use a different line element for this contribution, as we take into account the effect of the gravitational potential of the lens, for which cosmology is irrelevant. We rearrange Eq. (6) into
$k^{0}\equiv\frac{\mathrm{d}t}{\mathrm{d}\lambda}$ $=\pm\sqrt{-\frac{g_{ij}}{g_{00}}\frac{\mathrm{d}x^{i}}{\mathrm{d}\lambda}\frac{\mathrm{d}x^{j}}{\mathrm{d}\lambda}-\frac{m^{2}}{g_{00}}}$
@@
where the first term in the second bracket is the path length along the affine parameter, the contribution to the time delay of which we have calculated in detail in the previous section [see Sec. IV.1, Eq. (27)]. The second term in the second bracket is the Shapiro time delay, which arises because of the presence of the gravitational field [59]. This contribution depends on the path $\gamma(\lambda)$. The graviton mass acts as a correction to the Shapiro time delay.